# Reinforcing Question Answering Agents with Minimalist Policy Gradient Optimization

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable versatility, due to the lack of factual knowledge, their application to Question Answering (QA) tasks remains hindered by hallucination. While Retrieval-Augmented Generation mitigates these issues by integrating external knowledge, existing approaches rely heavily on in-context learning, whose performance is constrained by the fundamental reasoning capabilities of LLMs. In this paper, we propose Mujica, a Multi-hop Joint Intelligence for Complex Question Answering, comprising a planner that decomposes questions into a directed acyclic graph of subquestions and a worker that resolves questions via retrieval and reasoning. Additionally, we introduce MyGO (Minimalist policy Gradient Optimization), a novel reinforcement learning method that replaces traditional policy gradient updates with Maximum Likelihood Estimation (MLE) by sampling trajectories from an asymptotically optimal policy. MyGO eliminates the need for gradient rescaling and reference models, ensuring stable and efficient training. Empirical results across multiple datasets demonstrate the effectiveness of Mujica-MyGO in enhancing multi-hop QA performance for various LLMs, offering a scalable and resource-efficient solution for complex QA tasks.
@@ 1 Introduction
The advent of Large Language Models (LLMs) ouyang2022training ; achiam2023gpt has redefined expectations for Artificial Intelligence. Despite showing remarkable versatility, directly using LLMs to tackle Question Answering (QA) tasks remains challenging due to their lack of factual knowledge. Without sufficient information, LLMs may hallucinate and generate plausible answers, thereby undermining the reliability of QA systems. An intuitive solution involves injecting massive up-to-date factual knowledge into LLM’s internal parameters through continued pre-training or finetuning. However, such a process not only demands substantial computational resources, but also may result in the catastrophic forgetting problem kirkpatrick2017overcoming , undermining the models’ previously acquired knowledge. To address this issue, Retrieval-Augmented Generation (RAG) guu2020retrieval ; lewis2020retrieval , a technique enabling LLMs to incorporate external knowledge via prompts, has drawn significant attention among researchers.
To effectively address complex multi-hop questions, most advanced RAG systems PoG ; Search-o1 ; RoG ; shao2023enhancing ; khattab2022demonstrate adopt a divide-and-conquer strategy. This involves decomposing the question, retrieving information for each sub-question, and synthesizing a final answer. Such a system can be naturally conceptualized as an agent taking charge of making a plan on how to answer the question, interacting with external data sources, and conducting necessary reasoning ToG ; ToG2 ; Search-o1 . Since the backbone LLMs are not specifically pre-trained on these subtasks, existing approaches primarily rely on in-context demonstration as the sole guidance. However, in-context learning (ICL) ICL  ICL has several intrinsic limitations in the QA task. First, including few-shot demonstrations requires the LLM to process more tokens for every question, which not only reduces inference speed but also increases the risk of performance degradation due to attention dispersion caused by excessively sequence lengths. Second, for different types of complex questions, one may need to provide demonstrations tailored to the specific question type to ensure effectiveness. Furthermore, in-context learning operates as a form of analogy-based reasoning, which does not improve the underlying fundamental capability of LLMs to handle QA tasks. Consequently, the model may struggle with uncommon questions in specialized domains. Therefore, to develop a truly powerful and efficient QA agent, applying post-training on LLMs becomes essential.
@@ 2.1 Problem Definition
This work addresses the task of Open-Domain Multi-Hop Question Answering (MHQA). QA systems aim to provide correct answers to given questions. Compared with single-hop QA, MHQA necessitates reasoning across and aggregating information from multiple sources (hops) to derive the final answer mavi2024multi , which renders MHQA a particularly challenging task. We focus on a more general setting, the open-domain setting, in this work, where relevant information must be retrieved from large-scale corpora (e.g., knowledge graphs, text collections). Further details are provided in the Appendix.
The inherent complexity of MHQA, characterized by multi-stage planning requirements and iterative evidence synthesis, necessitates framing the solution mechanism as an autonomous agent. We define an agent as an entity capable of planning, taking actions based on the interactions with the environment, and engaging in sequential reasoning liu2025advances . We observe that LLMs equipped with RAG perform functions analogous to such an agent. Specifically, LLMs determine information requirements and search strategies (planning), formulate search queries (acting), and synthesize final answers from gathered evidence (reasoning). Recognizing this functional correspondence, we position the LLM itself as the core agent within our MHQA system.
@@ 3.3 Minimalist Policy Gradient Optimization (MyGO)
Given $\delta\in\mathbb{R}$, $\alpha\in R$, there exists $K\in\mathbb{R}^{+}$ such that the following inequality holds:
@@
$\mathrm{Var}(\pi^{>K})\propto 1/Z(\alpha)^{2}\;.$
@@ 4.1 Environment Design
To evaluate the effectiveness of our proposed agent workflow (Mujica) and training method (MyGO) on complex question answering tasks, we utilize four representative benchmark datasets: 2Wiki-MultihopQA (2Wiki) (ho2020constructing, ), QALD-10 (QALD-10, ), HotpotQA (yang2018hotpotqa, ), and MuSiQue (trivedi2021musique, ).
2Wiki and QALD-10 are KBQA benchmarks; HotpotQA and MuSiQue are text-based QA dataset. We use 2Wiki and HotpotQA for training, QALD-10 and MuSiQue for zero-shot evaluation. We assess QA performance using Exact Match (EM) and F1 score. For each dataset, these metrics are computed using the official evaluation scripts provided by the respective dataset creators. Comprehensive dataset statistics and precise definitions of the evaluation metrics are detailed in Appendix.

### v2

@@ Abstract
Large Language Models (LLMs) show remarkable versatility but face challenges in Question Answering (QA) due to knowledge cutoffs and hallucination. While Retrieval-Augmented Generation (RAG) helps by integrating external knowledge, current methods often depend on in-context learning, which can be limited by backbone models and prompt engineering. To overcome these limitations, we introduce Mujica (Multi-hop Joint Intelligence for Complex Question Answering), a novel multi-agent collaborative framework employing a divide-and-conquer strategy for QA. This framework is trained using MyGO (Minimalist policy Gradient Optimization), a new reinforcement learning method that directly samples and learns from the optimal policy. Empirical results across multiple datasets demonstrate that the Mujica-MyGO approach significantly enhances multi-hop QA performance for various LLMs.
@@ 1 Introduction
The advent of Large Language Models (LLMs) ouyang2022training ; achiam2023gpt has redefined expectations for Artificial Intelligence, offering new opportunities for Question Answering (QA) tasks due to their impressive language understanding and generation capabilities. However, directly applying LLMs to QA is challenging because of hallucination huang2025survey and knowledge cutoffs cheng2024dated . While injecting knowledge through continue learning is a possible solution, it is computationally expensive and risks catastrophic forgetting kirkpatrick2017overcoming , where models lose previously acquired knowledge. To address this issue, Retrieval-Augmented Generation (RAG) guu2020retrieval ; lewis2020retrieval , a technique enabling LLMs to incorporate external knowledge via prompts, has drawn significant attention among researchers.
Compared with single-hop questions which have one-turn retrieve-then-answer pipeline, most advanced RAG systems PoG ; Search-o1 ; RoG ; shao2023enhancing ; khattab2022demonstrate adopt a multi-turn pipeline to effectively address multi-hop questions. This involves decomposing the question, retrieving information for each sub-question, and synthesizing a final answer. Such a system can be naturally conceptualized as an agent taking charge of making a plan on how to answer the question, interacting with external data sources, and conducting necessary reasoning ToG ; ToG2 ; Search-o1 . However, existing approaches primarily rely on in-context demonstration as the sole guidance. However, in-context learning (ICL) ICL has several intrinsic limitations in the QA task. First, including few-shot demonstrations requires the LLM to process more tokens for every question, which not only reduces inference speed but also increases the risk of performance degradation due to attention dispersion caused by excessively sequence lengths. Second, for different types of complex questions, one may need to provide demonstrations tailored to the specific question type to ensure effectiveness. Furthermore, in-context learning operates as a form of analogy-based reasoning, which does not improve the underlying fundamental capability of LLMs to handle QA tasks. Consequently, the model may struggle with uncommon questions in specialized domains. Therefore, to develop a truly powerful and efficient QA agent, applying post-training on LLMs becomes essential.
@@ 2.1 Problem Definition
This work addresses the task of Open-Domain Multi-Hop Question Answering (MHQA). QA systems aim to provide correct answers to given questions. Compared with single-hop QA, MHQA necessitates reasoning across multiple sources (hops) mavi2024multi , which renders MHQA a more challenging task. We focus on the open-domain setting, where relevant information must be retrieved from large-scale corpora (knowledge graphs or text collections). Details are provided in the Appendix.
The inherent complexity of MHQA, characterized by multi-stage planning requirements and iterative evidence synthesis, necessitates framing the solution mechanism as an autonomous agent. We define an agent as an entity capable of planning, taking actions based on the environment, and engaging in sequential reasoning liu2025advances . We observe that LLMs equipped with RAG perform functions analogous to such an agent. Specifically, LLMs determine information requirements and search strategies (planning), formulate search queries (acting), and synthesize final answers from gathered evidence (reasoning). Recognizing this functional correspondence, we position the LLM itself as the core agent within our MHQA system.
@@ 3.3 Minimalist Policy Gradient Optimization (MyGO)
Given $\delta\in\mathbb{R}$, $\alpha\in\mathbb{R}$, there exists $K\in\mathbb{R}^{+}$ such that the following inequality holds:
@@
The following property holds when $\alpha\to 0$:
[…]
@@ 4.1 Environment Design
To evaluate the effectiveness of our proposed agent workflow (Mujica) and training method (MyGO) on complex question answering tasks, we utilize four representative benchmark datasets: 2Wiki-MultihopQA (2Wiki) (ho2020constructing, ), QALD-10 (QALD-10, ), HotpotQA (yang2018hotpotqa, ), and MuSiQue (trivedi2021musique, ). 2Wiki and QALD-10 are KBQA benchmarks; HotpotQA and MuSiQue are text-based QA dataset. We use 2Wiki and HotpotQA for training, QALD-10 and MuSiQue for zero-shot evaluation. We assess QA performance using Exact Match (EM) and F1 score. For each dataset, these metrics are computed using the official evaluation scripts provided by the respective dataset creators. Comprehensive dataset statistics and precise definitions of the evaluation metrics are detailed in Appendix.